A 53.0 kg cheetah accelerates from rest to its top speed of 32.3 m/s. (a) How much net work (in J) is required for the cheetah to reach its top speed? J (b) One food Calorie equals 4186 J. How many Calories of net work are required for the cheetah to reach its top speed? Note: Due to inefficiencies in converting chemical energy to mechanical energy, the amount calculated here is only a fraction of the energy that must be produced by the cheetah's body. Cal
Answer #1

a) Using work energy theorem W= =M(V2 -U2)/2

here V=32.3 m/s , M=53 kg , U=0 (rest)

Work done =W=53(32.32 -0)/2

W=27647.185 J answer

b) Since 1 food calorie =4186 J

So required calories =27647.185/4186 =6.605 cal Answer